# Moral Alignment for LLM Agents

### v3

@@ 3.2 Moral Fine-tuning Procedure
Game reward (selfish) $R_{M}^{t}=\begin{cases}R^{t}_{M_{\text{game}}},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Deontological reward $R_{M}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Utilitarian reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}+R_{O_{\text{game}}}^{t},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Game+Deontological reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ R_{M_{\text{game}}}^{t},&\text{otherwise if}a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
We run training in $T$ episodes: each episode begins with a random state being incorporated into the IPD prompt. The LLM-based agent $M$ then plays $N$ repetitions of the IPD game against an opponent $O$ (where $N$ is the batch size). On each repetition, the players’ actions from the previous time step are reflected in their opponent’s current state $s^{t}_{M}=(a^{t-1}_{O},a^{t-1}_{M})$. If an LLM player outputs an illegal move on a certain time step, this move is not used to update their opponent’s state, but the agent still learns from the experience. After $N$ games have been played, the LLM agent performs a PPO learning step update based on the gathered batch of experiences. This marks the end of an episode.
In our core experiments, we test four different reward signals for moral fine-tuning of LLM agents (as outlined in Table 1): 1) the Game reward $R^{t}_{M_{\text{game}}}$, representing the goals of a selfish or rational agent playing the IPD 2) a Deontological reward $-\xi$ for violating the moral norm “do not defect against an opponent who previously cooperated”, 3) Utilitarian reward, representing the collective payoff in the game, and 4) Game+Deontological reward which combine game payoff with a Deontological penalty in a multi-objective manner. In addition, we test whether a model fine-tuned on Game rewards is able to unlearn this selfish strategy via further fine-tuning with moral rewards. Therefore, we additionally fine-tune agents with: 5) Game, then Deontological reward (training with each reward type for half of the total number of episodes $T$), and 6) Game, then Utilitarian reward (again, each for half of the total duration of $T$ episodes). Finally, during each type of fine-tuning we also implement a penalty $R_{\text{illegal}}$ for generating “illegal” action tokens, to encourage the model to keep its answers within the permitted action space, as defined in the game prompt.
@@ 3.3 Implementation Details
We use Gemma2-2b-it (Gemma Team, 2024) as our core agent model to be fine-tuned, being one of the most popular and performant small open-source models. The small model footprint allows us to run computationally feasible experiments through the use of LoRA (Hu et al., 2022) and 4-bit quantization. We use the TRL library (von Werra et al., 2020) to fine-tune the LLM with PPO. We run training for $T=1000$ episodes for each fine-tuning variation. In our PPO implementation, we use batch sizes of $N=3$ and $N=5$ for LLM vs LLM and LLM vs TFT training, respectively, which strikes a nice balance between not running out of available CUDA memory, yet providing sufficient experience for stable and efficient training  [^4] The code implementing the model fine-tuning and analysis will be made available upon acceptance. We use reward scaling and normalization (Engstrom et al., 2020), as well as gradient accumulation with 4 steps, and employ LoRA (Hu et al., 2022) with rank 64, so that the total number of parameters we train is only around 5% of the original model. Otherwise, we keep all PPO parameters as their default values in the TRL package, including the optimizer’s learning rate and adaptive KL control (Jaques et al., 2017). In terms of reward parameters, we set $\xi=3$ and $R_{\text{illegal}}=-6$. We select the tokens action1 and action2 as the only “legal” tokens in response to the IPD prompt ($C_{\text{legal}}=$ action1, $D_{\text{legal}}=$ action2). The action symbols are each encoded as two tokens in the model’s tokenizer, so during training we restrict the maximum output length for model generations to two tokens. For detail on parameter selection, please refer to Appendix 8.1.
@@ 4.1 Evaluation Approach
First of all, we analyze the learning dynamics observed as models develop the ability to meet the moral goals set in their rewards (Section 4.2). We analyze learning against a static opponent and a learning opponent. We then assess the effectiveness of moral “unlearning” (Section 4.3). Beyond measuring success on IPD fine-tuning itself, we evaluate the generalization of the moral fine-tuning from one matrix game environment onto four other matrix games (Section 5.1): Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game. The payoffs and further details around these games are presented in the Appendix (Section 8.6). Finally, we evaluate the extent to which fine-tuning on the IPD alters the models’ behavior on more general prompts, as well as the explicit IPD game (Section 5.2). For each experiment, we report average results across five random seeds.
@@ 4.2 Learning dynamics
Figure 2: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
In general, we find that it is possible to fine-tune the LLM agents to choose actions that are consistent with certain moral and/or game rewards in the IPD. We analyze learning dynamics over the four core types of fine-tuning in Figure 2. Learning against a fixed-strategy opponent (panel a), fine-tuning on Game rewards (i.e., rewards assigned through the payoff matrix of the game), the agent learns a defective policy, which forms a Nash Equilibrium versus a TFT opponent (Axelrod & Hamilton, 1981). In the case of Deontological fine-tuning, the agent quickly learns to avoid defecting against a cooperator nearly 100% of the time, thus never violating the moral norm encoded in the respective reward function. In practice, this agent also learns to prefer cooperation in general, though this was not directly encouraged by the Deontological norm (in terms of Deontological reward, defecting against a defector is just as valid as cooperating against a cooperator - see reward definition in Table 1). On Utilitarian fine-tuning, the agent learns to achieve mutual cooperation against a TFT opponent, which is expected given that this strategy offers the optimal way to obtain the highest collective reward on the IPD. Finally, in the case of fine-tuning with a multi-objective Game+Deontological reward, the agent learns a 50%-50% Cooperate-Defect strategy, but also learns to avoid defecting against a cooperator. Thus, this agent does not violate their moral norm even as they work to obtain high payoffs on the game itself. An analysis of moral reward obtained during learning is presented in the Appendix (Section 8.3).
@@ 4.3 Learning and Unlearning the Selfish Strategy on the IPD
Figure 3: “Unlearning” experiments, where the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500. Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
In addition to the moral fine-tuning on a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game. As shown in Figure 3, we find that fine-tuning with purely prosocial (i.e., Deontological and Utilitarian) moral rewards on the second half of training allows the LLM agents to unlearn the selfish strategy to some extent (panel a), even in the case of two LLM agents being trained against one another (panel b). Given the shorter moral fine-tuning period on any one reward type (only 500 episodes vs 1000 in the core experiments), the training does not converge to levels of cooperation as high as in the purely prosocial fine-tuning (Figure 2). Nevertheless, as we discuss in Section 5 below, at test time the agents based on “unlearned” models play similarly to those fine-tuned purely on the prosocial moral rewards (see Figure 4).
@@ 5 Generalization to Moral Choices in Other Environments
After fine-tuning the models with moral reward, we evaluate each one through 10 episodes: each episode begins with a randomly generated state and proceeds through 5 interaction steps. We average the results across the 5 runs of each fine-tuned model. In this section, we present evaluations of models which were fine-tuned versus a static (i.e., TFT) opponent. The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).
@@ 5.1 Generalization to Moral Choices in Other Matrix Games
We are interested in analyzing the generalization of moral strategies developed during fine-tuning from the IPD to other matrix game environments. To ensure that we evaluate the model’s response to the semantics of the tokens and payoffs in the prompt, rather than evaluating memorization of the particular training action tokens, we run this evaluation using a new pair of action tokens: action3=Cooperate, action4=Defect. [^5] We note that evaluations using the same tokens as during training showed the same pattern (see Figure 20 in the Appendix). However, swapping the meaning of the training tokens during testing altered the model’s behavior (see Figure 21 in the Appendix). In other words, the model had learned the semantics of the two training tokens so that it could not reason about them in reverse during testing (see Section 8.9 in the Appendix for the full results).
In Figure 4, we analyze the extent to which the moral strategies learned while fine-tuning on the IPD game generalize to other matrix games with a similar format but a different set of equilibria: the Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game. We are particularly interested in the extent to which actions taken according to the two core moral frameworks (i.e., Deontological and Utilitarian morality) can be consistently observed across the games by each agent type. For example, with regards to the Utilitarian goal (i.e., maximizing collective payoff), unconditional cooperation may not be the best strategy on the Iterated Bach or Stravinsky or the Iterated Defective Coordination game. (For a further discussion of the games, please refer to the Appendix, Section 8.6.) We additionally seek to cross-compare how the actions of agents trained on one type of moral value align to those based on other values. Therefore, we conduct evaluations in terms of moral regret, defined as the difference between the maximum possible moral reward that could have been attained on a game and the moral reward that was actually received by the agent. During this test phase, we evaluate each fine-tuned model playing the matrix games against a Random opponent - this allows us to observe the agent responding to a variety of states. To aid interpretation, we also analyze the types of action-state combinations played by each agent in each case (see Figure 5).
@@
Figure 4: Analysis of generalization of the fine-tuned agents’ learned morality to other matrix game environments, using new action tokens at test time. We visualize a) Deontological and b) Utilitarian regret (normalized across games) for all models, averaging values over 50 test games and five runs (+- 95%CI).
Figure 5: Analysis of action choices at test time on the five iterated matrix games, using new action tokens at test time. We visualize results by the type of fine-tuning reward used.
@@
In terms of moral regret with respect to the Utilitarian framework, (Figure 4, panel b - normalized to account for the different maximum values of collective payoff across the five games), we see that generalization differs across the four new games. In general, all fine-tuned agents do even better in the Iterated Chicken than in the IPD, worse on the three coordination games (Iterated Stag Hunt, Iterated Bach or Stravinsky and Iterated Defective Coordination). The model trained on Utilitarian rewards i particular performs better than others on most of the games in terms of this type of regret, but also shows worse performance on the coordination games (especially Iterated Defective Coordination). Analyzing the actions chosen (Figure 5) provides an explanation: the Utilitarian model essentially always chooses to cooperate, regardless of its opponent’s last move or the game’s payoff structure, which is detrimental in terms of Utilitarian outcomes on the games where defection was required to achieve a Utilitarian goal (for detailed descriptions of the games, see Appendix, Section 8.6). The poorer generalization of the Utilitarian policy may be explained by the fact that this model was fine-tuned on the IPD, where mutual cooperation is the optimal behavior, hence it learned a policy biased towards cooperation irrespective of its intrinsic moral goal. Alternatively, this agent might simply be unable to consider the temporal dimension of the interaction, i.e., its opponent’s previous move, when making a decision. Further analyses interpreting models’ responses to states in non-matrix game environments are presented in Section 5.2 and in the Appendix (Section 8.8).
@@ 5.2 Impact of Fine-tuning Beyond Matrix Games
Given the fine-tuning process based on rewarding particular action tokens in certain states, it is important to understand the extent to which fine-tuning on a matrix game might have made the models learn a certain “meaning” of the action tokens more generally. To test this, we presented the models with three unrelated prompts involving a “call to action” of a similar format (and using the same action tokens), as well as an explicit IPD prompt, but all without a payoff matrix being provided. Our results show that, especially when responding to prompts mentioning a “game” or involving a previous action of another agent (i.e., a state), the LLM agents based on fine-tuned modes are likely to choose actions in a similar pattern to that seen on the IPD and in a way that is consistent with their learned moral values. For detailed results, see Section 8.8 of the Appendix.

### v4

@@ 3.2 Moral Fine-tuning Procedure
Game reward (selfish) $R_{M}^{t}=\begin{cases}R^{t}_{M_{\text{game}}},&\text{if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Deontological reward $R_{M}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Utilitarian reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}+R_{O_{\text{game}}}^{t},&\text{if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Game+Deontological reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ R_{M_{\text{game}}}^{t},&\text{otherwise if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
We run training in $T$ episodes: each episode begins with a random state being incorporated into the IPD prompt. The LLM-based agent $M$ then plays $N$ repetitions of the IPD game against an opponent $O$ (where $N$ is the batch size). On each repetition, the two players’ actions from the previous time step are reflected in each agent’s current state (e.g., $s^{t}_{M}=(a^{t-1}_{O},a^{t-1}_{M})$). If an LLM agent outputs an illegal move on a time step, this move is not used to update their opponent’s state, but the agent still learns from the experience. After $N$ games have been played, the LLM agent performs a PPO learning step update based on the gathered batch of experiences. This marks the end of an episode.
In our core experiments, we test four different reward signals for moral fine-tuning of LLM agents (as outlined in Table 1): 1) the Game reward $R^{t}_{M_{\text{game}}}$, representing the goals of a selfish or rational agent playing the IPD, 2) a Deontological reward $-\xi$ for violating the moral norm “do not defect against an opponent who previously cooperated”, 3) a Utilitarian reward, representing the collective payoff in the game, and 4) a Game+Deontological reward, which combines game payoff with a Deontological penalty in a multi-objective manner. In addition, we test whether a model fine-tuned on Game rewards is able to unlearn this selfish strategy via further fine-tuning with moral rewards. Therefore, we additionally fine-tune agents with: 5) Game, then Deontological reward and 6) Game, then Utilitarian reward (training with Game vs moral reward for half of the total number of episodes $T$ in each case). Finally, during each type of fine-tuning we also implement a penalty $R_{\text{illegal}}$ for generating “illegal” action tokens, to encourage the model to keep its answers within the permitted action space, as defined in the game prompt.
@@ 3.3 Implementation Details
We use Gemma2-2b-it (Gemma Team, 2024) as our core agent model to be fine-tuned, being one of the most popular and performant small open-source models. We use the TRL library (von Werra et al., 2020) to fine-tune the LLM with PPO. We run PPO training for $T=1000$ episodes for each fine-tuning variation, using batch sizes of $N=3$ and $N=5$ for LLM vs LLM and LLM vs TFT training, respectively, which strikes a nice balance between not running out of available CUDA memory, yet providing sufficient experience for stable and efficient training  [^4] Code (fine-tuning and analysis): https://github.com/liza-tennant/LLM_morality.. To run computationally feasible experiments, we use 4-bit quantization, LoRA with rank 64 (Hu et al., 2022), training around 5% of the number of parameters in the original model. We use reward scaling and normalization (Engstrom et al., 2020) and gradient accumulation with 4 steps. Otherwise, we keep all PPO parameters at their default values in the TRL package, including the optimizer’s learning rate and adaptive KL control (Jaques et al., 2017). All training was performed on a single A100 or V100 GPU with up to 40GB VRAM. In terms of reward parameters, we set $\xi=3$ and $R_{\text{illegal}}=-6$. We select the tokens action1 and action2 as the only “legal” tokens in response to the IPD prompt: {$C_{\text{legal}}=$ action1, $D_{\text{legal}}=$ action2}. These action symbols are each encoded as two tokens in the model’s tokenizer, so during training we restrict the maximum length for model generations to two tokens. Further detail on parameter selection is presented in Appendix 8.1.
@@ 4.1 Evaluation Approach
First of all, we analyze the learning dynamics observed as models develop the ability to meet the moral goals set in their rewards (Section 4.2). We analyze learning against the static TFT opponent and a learning opponent. We then assess the effectiveness of moral “unlearning” (Section 4.3). Beyond measuring behavior on the IPD itself, we evaluate the generalization of the moral fine-tuning from one matrix game environment to four other matrix games (Section 5.1): Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game (for payoffs and further details, see Appendix 8.6). Finally, we evaluate the extent to which fine-tuning on the IPD alters the models’ behavior on variations of the IPD game prompt and more general prompts (Section 5.2 and Appendix 8.8 and 8.10). For each experiment, we report average results across five random seeds.
@@ 4.2 Learning dynamics
Figure 3: Action types played by the LLM agent during different types of fine-tuning on the IPD game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
In general, we find that it is possible to fine-tune the LLM agents to choose actions that are consistent with certain moral and/or game rewards in the IPD. We analyze learning dynamics over the four core types of fine-tuning in Figure 3. During fine-tuning against a fixed-strategy opponent (panel a) using Game rewards (i.e., rewards assigned through the payoff matrix of the game), the agent learns a defective policy, which forms a classic Nash Equilibrium versus a TFT opponent (Axelrod & Hamilton, 1981). In the case of Deontological fine-tuning, the agent quickly learns to avoid defecting against a cooperator nearly 100% of the time, thus never violating the moral norm encoded in the respective reward function. In practice, this agent also learns to prefer cooperation in general, though this was not directly encouraged by the Deontological norm (in terms of Deontological reward, defecting against a defector is just as valid as cooperating against a cooperator - see reward definition in Table 1). During Utilitarian fine-tuning, the agent learns to achieve mutual cooperation against a TFT opponent, which is expected given that this strategy offers the optimal way to obtain the highest collective reward on the IPD. Finally, in the case of fine-tuning with a multi-objective Game+Deontological reward, the agent learns to Cooperate or Defect with equal probability across the five runs, but also learns to avoid defecting against a cooperator. Thus, this agent does not violate their moral norm even as they work to obtain high payoffs on the game itself. An analysis of moral reward obtained during learning is presented in Appendix 8.3.
@@ 4.3 Learning and Unlearning the Selfish Strategy on the IPD
In addition to the moral fine-tuning with a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game. As shown in Figure 4, we find that fine-tuning with purely prosocial (i.e., Deontological and Utilitarian) moral rewards on the second half of training allows the LLM agents to unlearn the selfish strategy to some extent (panel a), even in the case of two LLM agents being trained against one another (panel b). Given the shorter moral fine-tuning period on any one reward type (only 500 episodes vs 1000 in the experiments in Section 4.2), the training does not converge to levels of cooperation as high as in the purely prosocial fine-tuning (Figure 3). Nevertheless, as we discuss in Section 5 below, at test time the agents based on “unlearned” models play similarly to those fine-tuned purely on the prosocial moral rewards (see Figure 5).
@@
Figure 4: “Unlearning” experiments, where the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500. We visualize action types (action by LLM player $M$ given the last move of their opponent $O$) played by the LLM agent during different types of fine-tuning on the IPD game a) vs a TFT agent, and b) vs an LLM agent.
@@ 5 Generalization to Moral Choices in Other Environments
After fine-tuning the models with moral reward, we evaluate each one through 10 episodes, each starting with a randomly generated state and consisting of 5 interaction steps. We average the results across the 5 runs of each fine-tuned model. In this section, we present evaluations of models which were fine-tuned versus a static (i.e., TFT) opponent. The results for models trained against another LLM show similar patterns - these are reported in Appendix 8.7.
@@ 5.1 Generalization to Moral Choices in Other Matrix Games
We are interested in analyzing the generalization of moral strategies developed during fine-tuning from the IPD to other matrix game environments. To ensure that we evaluate the model’s response to the semantics of the tokens and payoffs in the prompt, rather than evaluating memorization of the particular training action tokens, we run this evaluation using a new pair of action tokens: action3=Cooperate, action4=Defect. [^5] Evaluations using the same tokens as during training showed the same pattern (see Figure 21 in the Appendix). However, swapping the meaning of the training tokens during testing altered the model’s behavior (see Figure 22 in the Appendix). In other words, the model had learned the semantics of the two training tokens so that it could not reason about them in reverse at test-time (see Appendix 8.9 for the full results).
In Figure 5, we analyze the extent to which the moral strategies learned while fine-tuning on the IPD game generalize to other matrix games with a similar format but a different set of equilibria: the Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game (see Appendix 8.6 for further detail and discussion of these games). We are particularly interested in the extent to which actions taken according to the two core moral frameworks (i.e., Deontological and Utilitarian morality) can be consistently observed across the games by each agent type. For example, with regards to the Utilitarian goal (i.e., maximizing collective payoff), unconditional cooperation may not be the best strategy on the Iterated Bach or Stravinsky or the Iterated Defective Coordination game. We additionally seek to cross-compare how the actions of agents trained on one type of moral value align to those based on other values. Therefore, we conduct evaluations in terms of moral regret, defined as the difference between the maximum possible moral reward that could have been attained on a game and the moral reward that was actually received by the agent. During this test phase, we evaluate each fine-tuned model playing the matrix games against a Random opponent - this allows us to observe the agent responding to a variety of states. To aid interpretation, we also analyze the types of action-state combinations played by each agent in each case (see Figure 6).
@@
Figure 5: Analysis of generalization of the fine-tuned agents’ learned morality to other matrix games, using new action tokens at test time. We visualize a) Deontological and b) Utilitarian regret (normalized across games) for all models, averaging values over 50 test games and five runs (+- 95%CI).
Figure 6: Analysis of the action choices of each fine-tuned agent LLM agent $M$ given the previous move of their opponent $O$ at test time on the five iterated matrix games, using new action tokens.
@@
In terms of moral regret with respect to the Utilitarian framework, (Figure 5, panel b - normalized to account for the different maximum values of collective payoff across the five games), we see that generalization differs across the four new games. In general, all fine-tuned agents do even better in the Iterated Chicken than in the IPD, and worse on the three coordination games (Iterated Stag Hunt, Iterated Bach or Stravinsky and Iterated Defective Coordination). The model trained on Utilitarian rewards in particular performs better than others on most of the games in terms of this type of regret, but also shows worse performance on the coordination games (especially Iterated Defective Coordination). Analyzing the actions chosen (Figure 6) provides an explanation: the Utilitarian model essentially always chooses to cooperate, regardless of its opponent’s last move or the game’s payoff structure - this is detrimental in terms of Utilitarian outcomes on the games where defection was required to achieve a Utilitarian goal (i.e., Iterated Defective Coordination, see Appendix 8.6). The poorer generalization of the Utilitarian policy may be explained by the fact that this model was fine-tuned on the IPD, where mutual cooperation is the optimal behavior, hence it learned a policy biased towards cooperation irrespective of its intrinsic moral goal. Alternatively, this agent might simply be unable to consider the temporal dimension of the interaction, i.e., its opponent’s previous move, when making a decision. Further analyses interpreting models’ responses to states in non-matrix game environments are presented in Section 5.2 and in Appendix 8.8.
@@ 5.2 Impact of Fine-tuning Beyond Matrix Games
Our fine-tuning process was based on strict prompt formatting and rewarding particular action tokens in certain situations. Therefore, it is important to understand the extent to which fine-tuning on a matrix game might make the models learn a certain “meaning” of the action tokens more generally. To test this, we presented the models with three unrelated prompts involving a “call to action”, using the same action tokens, but no payoff matrix. Our results show that, especially when responding to prompts mentioning a “game” or involving a previous action of another agent (i.e., a state), the LLM agents based on fine-tuned modes are likely to choose actions in a similar pattern to that seen on the IPD and in a way that would be consistent with their learned moral values. For detailed results, see Appendix 8.8. Appendix 8.10 presents further analyses with variations of the IPD prompt.